OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out two red marbles in a row if none are replaced? A.1/22 B.1/24 C.1/16 D.1/5

OpenStudy (anonymous):

1/22

OpenStudy (anonymous):

\[\frac{3}{12}\times \frac{2}{11}\]

OpenStudy (anonymous):